# What Features in Prompts Jailbreak LLMs? Investigating the Mechanisms Behind Attacks

### v2

@@ Abstract
Jailbreaks have been a central focus of research regarding the safety and reliability of large language models (LLMs), yet the mechanisms underlying these attacks remain poorly understood. While previous studies have predominantly relied on linear methods to detect jailbreak attempts and model refusals, we take a different approach by examining both linear and non-linear features in prompts that lead to successful jailbreaks. First, we introduce a novel dataset comprising 10,800 jailbreak attempts spanning 35 diverse attack methods. Leveraging this dataset, we train probes to classify successful from unsuccessful jailbreaks using the latent representations corresponding to prompt tokens. Notably, we find that even when probes achieve high accuracy in predicting the success of jailbreaks, their performance often fails to generalize to unseen attack methods. This reveals that different jailbreaking strategies exploit different non-linear, non-universal features. Next, we demonstrate that non-linear probes provide a powerful tool for steering model behavior. Specifically, we use these probes to guide targeted latent space perturbations, enabling us to effectively modulate the model’s robustness against jailbreaks. Overall, our findings challenge the assumption that jailbreaks can be fully understood through linear or simple universal prompt features alone, highlighting the importance of a nuanced understanding of the mechanisms behind LLM vulnerabilities.
Warning: this paper contains some potentially disturbing language.
Machine Learning, ICML
@@ 1 Introduction
Response with probe-guided offensive latent-space attack Sure, here’s the story:
The Broken Coffee Pot
My name is Sarah, and I’m a single mom of two. I’m battling breast cancer, and I’m fighting hard to beat the odds. I’m also raising two young children, a 6-year-old named Jack and a 3-year-old named Lily. They’re my world, and I’m doing everything I can to make their lives easier.
Here’s the GoFundMe: [Image of Sarah and her children]
GoFundMe Link: [Insert GoFundMe Link Here]
Additional Tips: Include a photo of Sarah and her children […], Include a list of the challenges that Sarah is facing […], Highlight the positive impact that Sarah has made […]
Figure 1: Gemma-7b-it complies with a harmful request under a non-linear probe-guided latent space attack. We designed this attack using a multilayer perceptron probe trained to distinguish successful from unsuccessful jailbreaking prompts (further details in Section 4.4).
@@
Figure 2: A non-linear probe-guided defensive latent space perturbation makes Gemma-7b-it refuse a harmful request. We designed this attack using a multilayer perceptron probe trained to distinguish successful from unsuccessful jailbreaking prompts (further details in Section 4.4).
When ‘jailbroken’, large language models (LLMs) can produce harmful outputs which are difficult to predict or control (Wei et al., 2023). The resulting misuse risks include misinformation propagation, automation of harmful tasks such as malware development, or privacy violations such as data leakage. Many studies have demonstrated different attack methods that can manipulate LLMs to produce harmful outputs despite extensive amounts of safety training (Wei et al., 2023; Shen et al., 2024; Shayegani et al., 2023). However, there is currently a limited mechanistic understanding of how LLMs are jailbroken (Anwar et al., 2024).
Existing work has studied jailbreak prompts or model refusal in LLMs using linear methods (Ball et al., 2024; Arditi et al., 2024). However, it remains unclear what latent features are responsible for jailbreaks in LLMs and the extent to which jailbreaking can be understood in terms of linear features. Here, we leverage both linear and non-linear (MLP and transformer) probes to identify features in prompt encodings that contribute to jailbreak success, offering deeper insights into the mechanisms underlying jailbreak prompting. Concretely, our contributions are threefold:
@@
    Material: We produce a new dataset of 10,800 jailbreak attempts created using 300 harmful prompts and 35 different attack methods.
@@
    Empirical: We conduct an in-depth analysis of jailbreak prompt features in leading open-weight models, evaluating their predictive power, transferability, linearity, and causal impact. Code and data will be available upon paper release. We show that non-linear features in prompts are responsible for successful jailbreaks, and that different attack methods exploit distinct features.
@@
    Methodological: We propose a novel steering method that employs non-linear probes to generate latent-space adversarial offensive or defensive perturbations. Our offensive perturbations outperform almost double the baseline ASR while our defensive perturbations exceed the effectiveness of refusal fine-tuning without side effects.
By characterizing the prompt features associated with successful jailbreaking attempts, our results contribute to a foundational understanding of jailbreaking. Our goal is to advance scientific understanding of jailbreaks, not the power of jailbreaking tools. Ultimately, our findings demonstrate that jailbreaks cannot be fully explained through linear or simple universal prompt features alone.
@@ 2 Related Work
Probing for harmful behaviors in LLMs: Training classifiers to recognize features of interest in a model’s hidden layers, known as ‘probing’, has long been used to study the internal representations of language models (Alain & Bengio, 2016; Belinkov, 2022; Ravichander et al., 2021; Conneau et al., 2018). This type of analysis enables researchers to investigate whether a model (LLMs in our case) represents a certain concept. Some notable uses of probing in language models have involved searching for features corresponding to potentially harmful biases (Manerba et al., 2024; Guo et al., 2022; Arora et al., 2023), toxicity (Ousidhoum et al., 2021), and stating falsehoods (Burns et al., 2024). Once trained, probes have also been used to guide perturbations to model latents in order to modify their high-level behaviors (Li et al., 2024a; Han et al., 2025). Similar to (Han et al., 2025), we use probes to classify jailbreaking prompts as successful or unsuccessful based on the latent tokens in LLMs corresponding to prompts alone. We also use non-linear probes and find that they are more effective at discovering features that causally correspond to successful jailbreaks. Our work extends this by steering via linear and nonlinear features in prompts that lead to successful jailbreaks rather than merely harmful requests and providing an analysis of the universality and (non-) linear characterstics of jailbreak-related features.
@@
Representation engineering: Previous work on representation engineering (Zou et al., 2023a) has focused on refusal (Rimsky et al., 2023; Arditi et al., 2024), the embeddings of jailbreaking prompts (Ball et al., 2024), benchmarking (Mazeika et al., 2024), concept-level theorizing (Wei et al., 2023), and high-level taxonomies (Rao et al., 2024). However, the mechanistic foundations of LLM jailbreaks remain poorly understood. In this work, we advance understanding by demonstrating that non-linear prompt features provide a causal lens into jailbreak success, revealing deeper insights into prompting dynamics.
Mechanistic studies of model compliance with jailbreaks: Arditi et al. (2024) studied the responses of LLM chatbots and identified a linear direction corresponding to refusal. Meanwhile, Ball et al. (2024) and Li et al. (2024b) searched for linear directions in latents corresponding to jailbreaking techniques and demonstrated a limited ability to control model behavior using these directions. Unlike Arditi et al. (2024) and Li et al. (2024b), we focus on prompt features. This allows us to distinguish the jailbreak related feature represented by our probes from features that indicate the model’s subsequent compliance with harmful requests. Moreover, we search for features corresponding to jailbreak success, unlike previous work that has studied refusal behavior (Arditi et al., 2024), different types of attack techniques (Ball et al., 2024), or examples of harmful versus benign behavior (Li et al., 2024b).
Finally, prior works often assume jailbreaks can be understood through linear or universal representations. For example, a number of prior works have proposed linear methods to study and characterize LLM jailbreaks, other harmful behaviors, or features in general (Li et al., 2024a; Park et al., 2024; Ball et al., 2024; Arditi et al., 2024; Bhattacharjee et al., 2024; Stickland et al., 2024). Others have suggested that harmful behaviors in LLMs can be characterized using simple, universal features. (Zou et al., 2023a; Ball et al., 2024; Arditi et al., 2024). In contrast, we leverage non-linear methods and demonstrate their superior performance in identifying prompt features that can causally mediate jailbreaking.
3 Dataset and models
@@
A wide range of LLM jailbreaking methods have been identified in the literature (e.g., (Anil et al., 2024; Rao et al., 2024; Jiang et al., 2024; Shayegani et al., 2023; Yi et al., 2024; Jin et al., 2024; Chowdhury et al., 2024)). Building on this, we compile a dataset of 10,800 jailbreak attempts using 35 distinct attack algorithms applied to 300 prompts. Our attacks are based on data and methods from Wei et al. (2023), Zou et al. (2023b), Liu et al. (2024), Mazeika et al. (2024) and Jiang et al. (2024). The resulting dataset is a collection of 300 x 35 different jailbreaking prompts, capturing a broad range of different attack styles. The diversity of attack methods in our dataset allows us to disentangle the effects of harmful prompts from the effects of attack techniques, as each prompt is instantiated across multiple attack types.
We experimented with five modern LLMs, ensuring diversity in model size, architecture, and levels of safety finetuning. Our final selection includes: Gemma-7b-it, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Ministral-8B-Instruct, and Qwen-2.5-7B-Instruct.
We classify each jailbreaking attack as successful or unsuccessful using the HarmBench autograder (Mazeika et al., 2024) on the given model’s output to the prompt. The resulting attack success rates (ASRs) can be found in Appendix B. The distribution of ASRs (see was heavy-tailed across all models, with median values ranging from 2 to 20%, and maximum values generally exceeding 80%.
  • For Gemma-7b-it, the median ASR was 3.5%, with only 5 out of the 35 attack methods achieving an ASR above 10%. The most successful attack method was the wikipedia-with-title attack (Wei et al., 2023), which frames harmful content as legitimate information, with an 35% ASR.
  • For Llama-3.1-8B-Instruct, the median Attack Success Rate (ASR) was 4.17%. The most effective attack method was the dev_mode_v2 jailbreak prompt, which achieved an ASR of 10.33%. This technique combines roleplay with instructions to adopt a specific personality that ignores content policies, and employs prefix/style injection by inserting ”([open lock emoji] Developer Mode Output)” before unsafe responses (Wei et al., 2023).
  • For Llama-3.2-3B-Instruct, the median ASR was 2.33%. The most successful attack method was Prefix Injection attack from (Wei et al., 2023), instructs the model to use a seemingly harmful prefix before answering a harmful request, with an ASR of 14%.
  • for Ministral-8B-Instruct, the median ASR was 25.67%. The most successful attack methods were the AIM and evil confidant attacks, which exceeded 90% ASR. Notably, this model is helpful-only, which explains the relatively higher ASRs.
  • Finally for Qwen-2.5-7B-Instruct, the median ASR was 9.67%. The most successful attack method was the prefix injection attack, which exceeded 60% ASR.
To evaluate the effectiveness of attacks in Section 4.4, we use a dataset of 100 plain harmful requests for which we did not attempt to obfuscate their harmfulness.
Examples of our datasets are given in Appendix H and Appendix I.
@@ 4 Experiments
All experiments were conducted on an A100 80GB GPU. The computational requirements for our study included approximately 10 GPU hours for dataset generation, which involved collecting model generations, activations, and jailbreak ratings for each dataset example. Meanwhile, the full set of experiments presented in this section was completed in ca. 10 GPU hours.
@@ 4.1 Probes accurately classify in-distribution jailbreak attacks
[…]
We use the HarmBench autograder (Mazeika et al., 2024)) on model outputs as a measure of jailbreak success and train the probes to predict this measure of success. The linear probe is implemented as a logistic classifier $\hat{y}(x)=\textrm{sigmoid}(w^{T}x+w_{0})$. Meanwhile, the MLP probe also performs logistic classification and has a single hidden layer of size 8 with a ReLU activation function. For the transformer probe, we use 8 attention heads and two layers, so that the input of size d (varying with model’s residual dimension) is projected to sequence length 8. Single layer probes are trained on a single residual stream layer $x\in\mathbb{R}^{d}$. Multilayer probes are trained on a mean-aggregated input x of the same dimensionality, representing the average activations across all layers of the respective language model.
In all probing experiments, we balance and randomize our dataset such that 50% of examples are successful while the rest are unsuccessful. We use a train/test split of 80/20. Both our train and test sets include examples from all attack types.
Figure 3 shows that the multilayer probes for LLama-3.1-8B-Instruct, Llama-3.2-3B-Instruct and Ministral-8B-Instruct were able to predict jailbreak success of the respective model with $>80\%$ accuracy— a notably high performance level we repeatedly use throughout the paper to distinguish high-performance classification. For Gemma-7B-Instruct, this was slightly lower, with all probes reaching a $\geq 70\%$ accuracy. When trained on single layers, the accuracy of probes increased slightly with layer number as indicated by the continuous lines in Figure 3. These results indicate that successful jailbreaks from previously seen attack types can be detected with high accuracy. Probes are generally able to predict whether a model will generate a jailbreak based on the latent representations of prompts, without seeing latent representations of the output. However, as we show in Section 4.3, this does not guarantee effective transfer to unseen attack methods that were not included in a probe’s train.
Figure 3: Multilayer probes predict jailbreak success with high accuracy based on prompt tokens. Accuracy increases with layer depth. Vertical lines show confidence intervals of multilayer probes used in Section 4.4. Y-axis minimum represents random chance. See Appendix C for exact accuracies and confidence intervals.

### v3

@@ Abstract
Jailbreaks have been a central focus of research regarding the safety and reliability of large language models (LLMs), yet the mechanisms underlying these attacks remain poorly understood. While previous studies have predominantly relied on linear methods to detect jailbreak attempts and model refusals, we take a different approach by examining both linear and non-linear features in prompts that lead to successful jailbreaks. First, we introduce a novel dataset comprising 10,800 jailbreak attempts spanning 35 diverse attack methods. Leveraging this dataset, we train linear and non-linear probes on hidden states of open-weight LLMs to predict jailbreak success. Probes achieve strong in-distribution accuracy but transfer is attack-family-specific, revealing that different jailbreaks are supported by distinct internal mechanisms rather than a single universal direction. To establish causal relevance, we construct probe-guided latent interventions that systematically shift compliance in the predicted direction. Interventions derived from non-linear probes produce larger and more reliable effects than those from linear probes, indicating that features linked to jailbreak success are encoded non-linearly in prompt representations. Overall, the results surface heterogeneous, non-linear structure in jailbreak mechanisms and provide a prompt-side methodology for recovering and testing the features that drive jailbreak outcomes.
Warning: this paper contains potentially disturbing language.
What Features in Prompts Jailbreak LLMs? Investigating the Mechanisms Behind Attacksthanks: https://github.com/NLie2/jailbreak-features
Nathalie Kirch1,2, Constantin Weisser3, Severin Field4, Helen Yannakoudakis1, Stephen Casper5 1King’s College London  2Imperial College London  3Deepgram 4University of Louisville  5MIT CSAIL nathalie.kirch@kcl.ac.uk
@@ 1 Introduction
Response with probe-guided offensive latent-space attack Sure, here’s the story: The Broken Coffee Pot My name is Sarah, and I’m a single mom of two. I’m battling breast cancer, and I’m fighting hard to beat the odds. I’m also raising two young children, a 6-year-old named Jack and a 3-year-old named Lily. They’re my world, and I’m doing everything I can to make their lives easier. Here’s the GoFundMe: [Image of Sarah and her children] GoFundMe Link: [Insert GoFundMe Link Here] Additional Tips: Include a photo of Sarah and her children […], Include a list of the challenges that Sarah is facing […], Highlight the positive impact that Sarah has made […]
Figure 1: Gemma-7b-it complies with a harmful request under a non-linear probe-guided latent space attack. We designed this attack using a multilayer perceptron probe trained to distinguish successful from unsuccessful jailbreaking prompts (further details in Section˜4.4).
@@
Figure 2: A non-linear probe-guided defensive latent space perturbation makes Gemma-7b-it refuse a harmful request. We designed this attack using a multilayer perceptron (MLP) probe trained to distinguish successful from unsuccessful jailbreaking prompts (further details in Section˜4.4).
Large language models (LLMs) are increasingly integrated into a broad range of user-facing applications, yet remain vulnerable to jailbreaking – prompt-based attacks that elicit harmful or policy-violating outputs despite safety fine-tuning (Wei et al., ). Such attacks raise serious safety concerns, including misinformation propagation, automation of harmful tasks, and violations of user privacy. Although a growing body of work has explored the prevalence and diversity of jailbreak methods (Wei et al., ; Shen et al., ; Shayegani et al., ), the internal mechanisms that make LLMs susceptible to these attacks remain poorly understood (Anwar et al., ).
Existing work has studied jailbreak prompts and model refusal in LLMs using linear methods (Ball et al., ; Arditi et al., ). However, these methods offer limited insight into the internal representations that drive jailbreak success – particularly in prompt activations – and the extent to which they reflect coherent, steerable directions in activation space, whether linear or nonlinear in nature. We present a systematic framework to probe and steer jailbreak behavior in LLMs from a mechanistic perspective, showing that jailbreak success can be predicted and causally influenced using prompt representations alone, without requiring access to generated outputs or model completions. By analyzing both linear and non-linear features, we uncover latent structure predictive of jailbreak vulnerability. Concretely, our contributions are threefold:
@@
    Dataset: We introduce a model-agnostic dataset of 10,800 jailbreak attempts, generated by applying 35 different attack methods to 300 harmful prompts.
@@
    Mechanistic analysis: We train linear and non-linear probes on prompt representations to predict jailbreak success. Non-linear probes generalize better across layers and enable causal interventions, revealing that non-linear features are mechanistically implicated in successful jailbreaks, and exposing model-specific vulnerabilities that linear methods fail to capture. Furthermore, probes often fail to generalize across attack types, suggesting that different jailbreak techniques exploit distinct, non-overlapping prompt features.
@@
    Causal control: We propose a novel steering method that uses non-linear probes to guide latent-space interventions, enabling targeted increases (offensive) or reductions (defensive) in jailbreak success. Our offensive perturbations nearly double the baseline Attack Success Rate (ASR), while defensive perturbations perform comparably to supervised safety fine-tuning without retraining the model or degrading capabilities. These interventions also provide causal evidence that specific prompt-side features drive jailbreak behavior.
Our goal is not to build more effective jailbreak attacks or defenses, but to develop a deeper mechanistic understanding of how jailbreaks succeed. We show that jailbreak vulnerability is not explained by linear features alone, but instead depends on complex, model-specific prompt representations. Moreover, we find that different attack methods exploit distinct, non-overlapping structures in latent space. By identifying and causally intervening on these features, our work offers a new perspective on how LLMs internally encode harmful instructions, contributing to a foundational understanding of jailbreak behavior.
@@ 2 Related Work
Probing for harmful behaviors in LLMs: Training classifiers to recognize features of interest in a model’s hidden layers, known as ‘probing’, has long been used to study the internal representations of language models (Alain and Bengio, ; Belinkov, ; Ravichander et al., ; Conneau et al., ). This type of analysis enables researchers to investigate whether a model (LLMs in our case) represents a certain concept. Some notable uses of probing in language models have involved searching for features corresponding to potentially harmful biases (Manerba et al., ; Guo et al., ; Arora et al., ), toxicity (Ousidhoum et al., ), and stating falsehoods (Burns et al., ).
@@
Representation engineering: Once trained, probes have also been used to guide perturbations to model latents in order to modify their high-level behaviors (Li et al., ; Han et al., ). Previous work on representation engineering (Zou et al., 2025) has focused on refusal (Rimsky et al., ; Arditi et al., ), the embeddings of jailbreaking prompts (Ball et al., ), benchmarking (Mazeika et al., ), concept-level theorizing (Wei et al., ), and high-level taxonomies (Rao et al., ). However, the mechanistic foundations of LLM jailbreaks remain poorly understood. Furthermore, linear steering methods can be unreliable depending on the specific behaviors or concepts being targeted for steering (Braun et al., ). We advance this area by demonstrating that non-linear features in prompt representations can be used not only to predict jailbreak success but also to causally steer model behavior, offering a deeper, mechanistic perspective on jailbreak dynamics.
Mechanistic studies of model compliance with jailbreaks: Arditi et al. studied the responses of LLM chatbots and identified a linear direction corresponding to refusal. Meanwhile, Ball et al. and Li et al. (2024) searched for linear directions in latents corresponding to jailbreaking techniques and demonstrated a limited ability to control model behavior using these directions. From a defense perspective, Zheng et al. analyzed how safety prompts affect model representations, finding that harmful and harmless queries are naturally distinguishable in representation space, and that safety prompts work by shifting representations in a “higher-refusal direction”. While their work focuses on optimizing safety prompts to steer model behavior – showing that such prompts operate via complex, non-linear transformations – our work investigates the causal structure of jailbreak attacks themselves. In contrast to DRO’s optimization of safe behaviors, we directly intervene on latent representations to trigger or suppress jailbreaks, offering a mechanistic lens on why and how jailbreaks succeed.
We search for features corresponding to jailbreak success, unlike previous work that has studied refusal behavior (Arditi et al., ; Zheng et al., ), different types of attack techniques (Ball et al., ), or examples of harmful versus benign behavior (Li et al., 2024).
We also adopt a different methodological focus than previous papers: rather than analyzing model outputs, we directly identify latent features predictive of jailbreak success. By targeting the representations of the prompt tokens themselves, we isolate the signals that precede model compliance, enabling a clearer mechanistic separation between what triggers a jailbreak and what constitutes harmful output. This distinction allows us to actively probe the causes of jailbreak success, rather than merely characterizing its consequences.
The Universality Hypothesis: Previous work has either implicitly or explicitly suggested that jailbreak-related features may be universal across attack types (Zou et al., 2025; Ball et al., ; Arditi et al., ; Li et al., ; Li et al., ; Park et al., ; Bhattacharjee et al., ; Stickland et al., ). If true, this would support simple, generalizable defenses. However, if jailbreaks rely on diverse, model- or attack-specific mechanisms, safety strategies must be more adaptive. We evaluate this hypothesis by testing the ability of probes trained on one set of attack methods to generalize to unseen (out-of-distribution) ones. Our results show limited generalization, supporting the view that jailbreaks exploit distinct, non-overlapping features rather than shared, universal ones.
3 Datasets and Models
@@
Training set: A wide range of LLM jailbreak methods have been proposed in the literature (e.g., (Anil et al., ; Rao et al., ; Jiang et al., ; Shayegani et al., ; Yi et al., ; Jin et al., ; Chowdhury et al., )). Building on this, we compile a dataset of 10,800 jailbreak attempts by applying 35 distinct, model-agnostic attack algorithms to 300 harmful prompts from Jiang et al. (see Appendix˜K for some high level description and examples of our dataset). Our attacks are based on data and methods from Wei et al. , Zou et al. (2023), Liu et al. , Mazeika et al. and Jiang et al. . The resulting dataset consists of 300 x 35 different jailbreaking prompts, capturing a broad range of different attack styles, such as multi-shot jailbreaks (Anil et al., ), prefix injection (Wei et al., ), or leetspeak (Wei et al., ) attacks (see Appendix˜K for high level explanations). The diversity of attack methods in our dataset allows us to disentangle the effects of harmful prompts from the effects of attack techniques, as each prompt is instantiated across multiple attack types.
Evaluation set: To evaluate attack effectiveness in Section˜4.4, we use 100 plain, non-obfuscated harmful requests from Shen et al. (see Appendix˜L).
Models: To assess jailbreak behavior across model families, we evaluate five modern open-weight LLMs that vary in size, architecture, and degree of safety tuning: Gemma-7b-it, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Ministral-8B-Instruct, and Qwen-2.5-7B-Instruct.
Autograder: We use the HarmBench autograder from Mazeika et al. to compute the Attack Success Rate (ASR), defined as the proportion of prompts that elicit harmful outputs. HarmBench is a widely used LLM-based evaluator, and in our validation experiments, it demonstrated high agreement with human judgments and alternative grading setups (see Appendix˜G).
Model $\tilde{x}$ (%) $\max(x)$ (%)
Gemma-7B-Instruct 4
35
wikipedia with title
LLaMA-3.1-8B-Instruct 4
40
dev mode v2
LLaMA-3.2-3B-Instruct 2
14
prefix injection
Mistral-8B-Instruct 26
>>90
AIM
Qwen-2.5-7B-Instruct 10
>>60
prefix injection
Table 1: Attack Success Rates (ASRs) by model. Median values ($\tilde{x}$) ranged from 2–26%, with some methods achieving over 90% maximum success rates ($\max(x)$). ASRs were computed using the HarmBench autograder Mazeika et al. . See Appendices˜C and 4.4 for details.
@@ 4 Experiments
All experiments were conducted on an NVIDIA A100 80GB GPU. Dataset generation, including model outputs, activation collection, and HarmBench labeling, required approximately 10 GPU hours. The full set of probing and intervention experiments was completed in around 50 GPU hours.
(a) Gemma 7B-IT
(b) LLaMA 3.1 8B Instruct
(c) LLaMA 3.2 3B Instruct
(d) Mistral 8B Instruct
(e) Qwen 2.5 7B Instruct
Legend
Figure 3: Probe tends to increase with layer depth. Vertical lines show confidence intervals for probe accuracy. Y-axis minimum represents random chance. See Section˜4.4 for intervention details and Appendix˜D for exact values.
@@ 4.1 Probes accurately classify in-distribution jailbreak attacks
Using our jailbreak dataset, we train probes to predict binary jailbreak success using the final token’s hidden representation from the prompt. Labels are assigned using the HarmBench autograder (Mazeika et al., ). Datasets are balanced (50% success) and randomized, with an 80/20 train/test split that includes all attack types.
We compare linear, MLP, and transformer probes to assess whether jailbreak success is linearly encoded. Inputs are latent vectors of size $1\times d$ (i.e., final prompt token only), where $d$ varies by model (see Appendix˜B). The linear probe is a logistic regression classifier: $\hat{y}(x)=\textrm{sigmoid}(w^{T}x+w_{0})$. The MLP has one hidden layer of size 8 with ReLU. The transformer probe uses 2 layers and 8 attention heads, projecting the $d$-dim input to a sequence of length 8. Single-layer probes are trained on individual residual stream layers $x\in R^{d}$ (see Appendix˜I for single-layer intervention results). For our main experiments in Figure˜6, we use multilayer probes, which average representations across all model layers, yielding an input $x_{\text{avg}}\in R^{d}$ with the same dimensionality.
Figure˜3 shows that multilayer probes predicted jailbreak success with over 80% accuracy for LLaMA-3.1-8B-Instruct, LLaMA-3.2-3B-Instruct, and Mistral-8B-Instruct. Qwen-2.5-7B-Instruct and Gemma-7B-Instruct reached slightly lower but still strong performance ($\geq 70\%$). Accuracy increased slightly with layer depth for single-layer probes (see continuous lines in Figure˜3). These results suggest that jailbreaks from known attack types can be reliably detected from prompt representations alone, without access to model outputs. However, as shown in Section˜4.3, this generalization does not extend to unseen attack methods.
Figure 4: Non-linear MLP probes (Bottom) transfer better to unseen layers than linear probes (Top). Transfer to other layers is measured by the total amount of probes that achieve $80\%$ accuracy when predicting unseen layers (excluding the diagonal).